Directions: Evaluate the following functions.



1. F(x) =x - x² find f(2)



2. g(a) = 3a - 2 find g(1)



3. g(x)=x²+7 find g(-2)



4.g(x)=3x-3 find g(-6)



5.h(x) = 4x-2 find h(3)

1
Expert's answer
2022-03-07T23:37:05-0500

1) "F(2)=2-2^2=-2."

2) "g(1)=3-2=1."

3) "g(-2)=(-2)^2+7=11."

4) "g(-6)=3(-6)-3=-21."

5) "h(3)=4*3-2=10."